# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v3

@@ Abstract
Large Language Models (LLMs) offer powerful capabilities in code generation, natural language understanding, and domain-specific reasoning. Their application to quantum software development remains limited, in part because of the lack of high-quality datasets both for LLM training and as dependable knowledge sources. To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories. Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline. Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness. Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
Index Terms:
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
@@ I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16].
@@
    RAG-Based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance. Evaluated GPT-4o mini, Claude 3.5, Qwen 7B, and LLaMa 4 Maverick 17B under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
Summary of Key Results: Retrieval augmentation with PennyLang yields substantial relative gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66% ($\approx$32.95% relative increase), and LLaMa 4 improves from 78.78% to 84.84% ($\approx$7.7% relative increase). In contrast, stronger baselines (GPT-4o-mini and Claude 3.5 Sonnet) do not benefit from naive full-context retrieval, their best performance is achieved without retrieval (85.60% and 95.07%, respectively), with full retrieval resulting in slight degradation. Detailed analysis is presented in Section 5.
@@ II Background and Related Work
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [40] and automated code generation [41], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows. This section presents an in-depth review of related work in three key areas:
@@ II-A LLMs for Code Generation in Classical Programming
II-B LLMs for Quantum Programming
Figure 2: Methodology for PennyLang Quantum Code Generation and Evaluation. We collect and refine quantum code from GitHub, textbooks, and PennyLane documentation, then use GPT-4o to convert cleaned snippets into instruction–query pairs (with code and tests). These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts. GPT-4o-mini, Claude 3.5 Sonnet, LLaMa 4-17B Maverick, and Qwen 2.5-7B-Instruct-Turbo are evaluated on functional correctness to quantify the benefit of PennyLang dataset.
@@ III-A1 Data Collection from GitHub Repositories
  • Non-PennyLane files were filtered by checking for explicit import pennylane as qml statements.
@@
Since PennyLane maintains backward compatibility and rarely deprecates features [19], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code. This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,321 samples were collected from the official PennyLaneAI GitHub repository [39].
@@ III-A2 Extracting Quantum Code from Books
    Manually verification with explanations from the books, converting descriptive text into structured comments.
@@ III-A4 Dataset Summary and Statistics
The final dataset containing 3,347 PennyLane code samples is shown in Table III.
@@
Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository + others) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
III-B Tokenization and Formatting for LLM Training
This section details the preprocessing pipeline for preparing our dataset for (LLM) training/fine-tuning, focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Handling Variable-Length Code Samples
The collected PennyLane code exhibit varying lengths after tokenization. Since transformer-based models require fixed-length inputs, we employ padding to standardize sequence lengths, ensuring that all samples match the longest sequence length $N$. Following standard practice, left padding was applied to align shorter samples to $N$ tokens, ensuring batch processing efficiency and minimizing computational overhead.
III-B2 Padding Strategies and Tokenizer Variations
@@
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like Qwen 2.5 Coder, GPT-4o Mini, Claude 3.5 Sonnet, and LLaMa 4. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
III-B3 Applying Attention Masking
To improve transformer efficiency, an attention mask [51] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table IV shows the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during training.
@@ III-B3 Applying Attention Masking
  • Impact on Training: The self-attention mechanism [50] ensures that padding tokens do not contribute to gradients, thereby reducing unnecessary computations and preventing misleading dependencies.
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To improve PennyLane code generation by reducing hallucinations and outdated syntax—and to better align outputs with PennyLane best practices—we adopt a retrieval-augmented generation (RAG) [52] approach. We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction. The PennyLang dataset, formatted as instruction–response pairs, is embedded using OpenAIEmbeddings and stored in a Chroma vector database for efficient similarity search. At inference, user queries are embedded, relevant context is retrieved, and the combined prompt is sent to the target LLM. Table V summarizes the key PennyLang RAG framework components, including their role in dataset loading, embedding storage, prompt formatting, retrieval chaining, and interfacing with multiple LLM backends.
@@
PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o-mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
@@ IV-A Response Lengths Across Feature Groups
Figure 3 provides a box plot representation of response lengths across different feature groups.
@@ IV-B Instruction and Response Length Distributions
Figure 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
@@ IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Figure 5. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
@@
TABLE VI: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o-mini, and Claude 3.5 Sonnet on PennyLane
Metric Qwen 7B LLaMa 4 GPT-4o-mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Total Notebooks 1320 1320 1320 1320 Working Notebooks 467 495 480 95 901 875 895 799 712 780 755 625 899 911 856 1001 Success Rate (%) 41.66 45.07 43.18 8.71 84.84 84.84 82.17 78.78 80.30 84.84 81.81 85.60 89.01 92.44 87.50 95.07 Pass@1 87 100 88 19 179 171 179 150 144 158 149 134 187 182 170 194 Pass@2 99 109 103 21 198 202 199 178 171 189 175 175 210 218 203 230 Pass@3 104 113 109 21 212 214 209 191 190 207 195 199 223 232 214 238 Pass@4 109 117 110 22 216 217 212 202 202 219 209 215 230 242 230 246 Pass@5 110 119 114 23 224 224 217 208 212 224 216 226 235 244 231 251
@@ IV-D Feature Category Association with Quantum Functions
To analyze how different quantum functionalities relate to feature categories, we examine the confusion matrix in Figure 6.
Figure 6: Confusion matrix showing the relationship between newly defined semantic categories and various PennyLane quantum features. The color intensity represents the frequency of occurrences, where darker shades indicate higher associations between a given category (rows) and a specific quantum feature (columns). This visualization provides insights into how different quantum functionalities are distributed across conceptual categories, aiding dataset structuring and LLM-based learning tasks.
@@ IV-E Insights and Implications
Findings from Figures 4, 3, 6, and 5 offer several insights:
@@ V Evaluation
To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts. Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs. The evaluation utilized the Pass@5 metric, where each model generates five candidate solutions per test, and success rate (%) is computed as $\text{pass@5}/{264}\times 100$, where pass@5 indicates the number of test cases for which at least one correct solution was generated among the five attempts. In total, 1320 notebooks were generated and executed per model (264 tests × 5 completions), ensuring a rigorous, code-level assessment of correctness. Table VII lists the details for evaluated models.
TABLE VII: LLMs Used in the RAG-Based Evaluation Framework
@@
Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)    $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)    $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Code generation (planned evaluation) GPT-4o Mini (OpenAI)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)    $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer    $\bullet$ Usage: Alternative model for response generation
@@
Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4. These models showed clear improvements when used with our dataset. Notably, RAG with 75% context coverage often outperformed the full-context setting. This suggests that a moderate amount of relevant context may reduce noise and improve grounding, while excessive retrieval might introduce redundancy or irrelevant tokens.
@@
For comparison, we also tested GPT-4o-mini and Claude 3.5. As expected, these commercial models demonstrated high performance even without RAG. However, the added context from our dataset provided limited benefit. This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.

### v4

# A PennyLane-Centric Dataset to Enhance LLM-based Quantum Code Generation using RAG

@@ Abstract
Large Language Models (LLMs) offer powerful capabilities in code generation, natural language understanding, and domain-specific reasoning. Their application to quantum software development remains limited, in part because of the lack of high-quality datasets both for LLM training and as dependable knowledge sources. To bridge this gap, we introduce PennyLang, an off-the-shelf, high-quality dataset of 3,347 PennyLane-specific quantum code samples with contextual descriptions, curated from textbooks, official documentation, and open-source repositories. Our contributions are threefold: (1) the creation and open-source release of PennyLang, a purpose-built dataset for quantum programming with PennyLane; (2) a framework for automated quantum code dataset construction that systematizes curation, annotation, and formatting to maximize downstream LLM usability; and (3) a baseline evaluation of the dataset across multiple open-source and commercial models, including ablation studies, all conducted within a retrieval-augmented generation (RAG) pipeline. Using PennyLang with RAG substantially improves performance: for example, Qwen 7B’s success rate rises from 8.7% without retrieval to 41.7% with full-context augmentation, and LLaMa 4 improves from 78.8% to 84.8%, while also reducing hallucinations and enhancing quantum code correctness. Moving beyond Qiskit-focused studies, we bring LLM-based tools and reproducible methods to PennyLane for advancing AI-assisted quantum development.
@@ I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [15, 46, 26]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [2], where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer, highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [39, 56, 23, 36, 21, 57, 24, 28, 42, 13, 22, 30, 37, 29, 20].
@@
    RAG-based Evaluation Framework: Designed a RAG evaluation framework to assess the dataset’s impact on model performance. Framework was evaluated on seven LLMs, including both open-source and commercial models under both retrieval-augmented and non-augmented settings, scoring outputs on functional correctness.
Summary of Key Results: Retrieval augmentation with PennyLang yields substantial gains for smaller open-source models: Qwen 7B’s success rate rises from 8.71% to 41.66%, and LLaMa 4 improves from 78.78% to 84.84%. In contrast, stronger baselines (commercial models) do not benefit significantly from naive full-context retrieval. For example, the best performance is achieved by Claude 3.5 Sonnet without retrieval (95.07%), with full retrieval resulting in slight degradation. Detailed analysis is presented in Section V.
@@ II Background and Related Work
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [50] and automated code generation [25], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [12], with little attention to the PennyLane framework [7, 6, 5], despite its popularity in hybrid quantum-classical workflows. This section presents an in-depth review of related work in three key areas:
@@ II-A LLMs for Code Generation in Classical Programming
II-B LLMs for Quantum Programming
Figure 2: Methodology for PennyLang Quantum Code Generation and Evaluation. We collect and refine quantum code from GitHub, textbooks, and PennyLane documentation, then use GPT-4o to convert cleaned snippets into instruction–query pairs (with code and tests). These examples are embedded in Chroma, and LangChain performs MMR-based retrieval to produce RAG-augmented versus vanilla prompts. Both open-source and commercial models are evaluated on functional correctness to quantify the benefit of the PennyLang dataset.
@@ III-A1 Data Collection from GitHub Repositories
Since PennyLane maintains backward compatibility and rarely deprecates features [7], we did not impose a time constraint on data collection, ensuring a broad dataset scope. Additionally, introductory comment blocks containing license information, author details, or unrelated metadata were removed to retain only executable and relevant code. This curation process resulted in a high-fidelity dataset of PennyLane-specific implementations. In total, 1,3211{,}321 samples were collected from the official PennyLaneAI GitHub repository [43].
@@ III-A2 Extracting Quantum Code from Books
    Manual verification with explanations from the books, converting descriptive text into structured comments.
@@ III-A4 Dataset Summary and Statistics
The final dataset containing 3,3473{,}347 PennyLane code samples is shown in Table III.
@@
Source Samples Collected Unofficial GitHub repositories 1,952 GitHub (PennyLaneAI repository) 1,321 Official PennyLane Documentation 53 Quantum Computing Books 21 Total Dataset Size 3,347
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
III-B Tokenization and Formatting for LLM
This section details the preprocessing pipeline for preparing our dataset for (LLM), focusing on tokenization, padding strategies, and attention masking to optimize input handling within the transformer architecture.
III-B1 Padding Strategies and Tokenizer Variations
@@
In our pipeline, we employ left padding as it aligns with the causal decoder-only transformer architectures used in models like GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini. This ensures that masked self-attention restricts predictions to depend only on preceding tokens, maintaining auto-regressive consistency.
III-B2 Applying Attention Masking
To improve transformer efficiency, an attention mask [4] was applied post-padding. This mask helps the model distinguish between valid input tokens and padding tokens, ensuring that computation is focused on meaningful data. Table IV shows the attention mask matrix applied to a batch of tokenized samples. Here, the "<pad>" tokens are assigned 0s, instructing the model to ignore them during processing.
@@ III-B2 Applying Attention Masking
  • Impact: Self-attention [52] ensures that padding tokens do not influence attention computations, thereby reducing unnecessary computations and preventing misleading dependencies.
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To improve PennyLane code generation by reducing hallucinations and outdated syntax, and to better align outputs with PennyLane best practices, we adopt a retrieval-augmented generation (RAG) [31] approach. We implement the pipeline in LangChain with modular components for document ingestion, vector embedding, retrieval, and prompt construction. The PennyLang dataset, formatted as instruction–response pairs, is embedded using OpenAIEmbeddings and stored in a Chroma vector database for efficient similarity search. At inference, user queries are embedded, relevant context is retrieved, and the combined prompt is sent to the target LLM. Table V summarizes the key PennyLang RAG framework components, including their role in dataset loading, embedding storage, prompt formatting, retrieval chaining, and interfacing with multiple LLM backends.
@@
PennyLang RAG Framework Components DirectoryLoader (Document Loader)    $\triangleright$ Loads PennyLane dataset samples as text files from directory Chroma (Vector Database)    $\triangleright$ Stores embeddings for retrieval using similarity search OpenAIEmbeddings (Embedding Model)    $\triangleright$ Converts text into vector embeddings for indexing and retrieval ChatPromptTemplate (Prompt Formatting)    $\triangleright$ Defines structured prompts for AI models to ensure consistency create_retrieval_chain (Retrieval Pipeline)    $\triangleright$ Creates pipeline integrating stored embeddings for improved responses ChatOpenAI (OpenAI LLM Interface)    $\triangleright$ Connects to OpenAI models (GPT-4o Mini) for response generation ChatAnthropic (Anthropic LLM Interface)    $\triangleright$ Connects to Anthropic Claude models (Claude 3.5 Sonnet) for responses
@@ IV-A Response Lengths Across Feature Groups
Fig. 3 provides a box plot representation of response lengths across different feature groups.
@@ IV-B Instruction and Response Length Distributions
Fig. 4 illustrates the distribution of instruction and response lengths, highlighting key patterns in the dataset. Instruction lengths, as shown in left plot, follow a normal distribution with a peak around 35 words. The majority of instructions fall within the 30–40 word range, while longer prompts exceeding 60 words are relatively rare. In contrast, response lengths exhibit a right-skewed distribution, as shown in the middle plot. Most responses are under 2500 words, with a distinct peak around 1000 words, though some responses exceed 10,000 words, indicating instances of verbose outputs. The instruction-response correlation, illustrated in the rightmost plot, reveals a positive correlation between instruction and response length. While longer instructions tend to produce longer responses, variability increases significantly beyond 50 words, highlighting the influence of instruction complexity on response verbosity.
@@ IV-C Feature Group Usage Across Quantum Domains
The relationship between feature groups and quantum computing domains, such as machine learning, hardware, chemistry, and optimization, is illustrated in Fig. 5. Quantum Machine Learning relies predominantly on quantum gates and operators, alongside error handling techniques, reflecting its dependence on structured quantum transformations and noise mitigation. In contrast, Quantum Chemistry and Optimization exhibit relatively lower usage of mathematical functions, suggesting that these domains rely on alternative computational approaches rather than extensive arithmetic operations. Quantum Hardware applications, however, demonstrate a balanced distribution across various feature groups, indicating a broader, more integrated usage of quantum functionalities across different operational layers.
@@ IV-E Insights and Implications
Findings from Figures 3, 4, and 5 offer several insights:
@@ V Evaluation
To assess the effectiveness of our proposed dataset for quantum code generation, we designed a benchmark suite comprising 264 test cases, each tailored to evaluate various PennyLane functionalities and quantum programming concepts. The test cases (prompts) cover a broad spectrum of tasks, ranging from basic circuit construction to advanced hybrid algorithm design, encompassing state preparation, parameterized gate implementation, differentiation routines, and optimization workflows. Several tasks address realistic applications, including molecular energy estimation via the Variational Quantum Eigensolver (VQE) for systems such as $\mathrm{H_{2}}$ and $\mathrm{LiH}$, as well as hybrid quantum–classical models trained on toy datasets for classification and regression. An example of an advanced prompt is:
“Write a PennyLane program that implements a Variational Quantum Eigensolver to estimate the ground-state energy of the $\mathrm{H_{2}}$ molecule using the default.qubit device. Define the molecular Hamiltonian, construct a parameterized ansatz circuit, and optimize its parameters using gradient-based optimization until convergence.”
Each test was used as a prompt within an RAG pipeline and evaluated across both open-source and commercial LLMs under different context settings (100%, 75%, 50%, and 0%), where 100% denotes full retrieved context and 0% corresponds to no RAG. This setup was used to evaluate the effect of retrieval on code generation performance. The evaluation employed pass@1 through pass@5, where each model generated five candidate solutions per test. The success rate (%) was computed as $\texttt{pass@k}/264\times 100$, where pass@k denotes the number of test cases, out of 264, for which at least one correct solution was obtained within the first k attempts. A generated solution was counted as correct if the notebook executed successfully and satisfied the corresponding test-case checks for the given inputs. The evaluation focused on whether the generated code correctly implemented the required task and produced the expected valid output for that test case. For tasks involving learning or optimization, a solution was considered correct if the requested workflow executed properly and returned the required type of result, regardless of the final performance value. In total, 1,320 notebooks were generated and executed per model (264 tests $\times$ 5 completions), providing a rigorous code-level functional evaluation. Table VI lists the evaluated models.
TABLE VI: LLMs Used in the RAG-based Evaluation Framework
@@
Large Language Models in RAG Pipeline Qwen2.5-7B-Instruct-Turbo (Alibaba Cloud)     $\bullet$ Parameters: 7B $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Generates PennyLane code LLaMa 4 Maverick (Meta)     $\bullet$ Parameters: 17B $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Code generation GPT-4o Mini (OpenAI)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Evaluates and scores generated responses Claude 3.5 Sonnet (Anthropic)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Alternative model for response generation Gemini 2.5 Flash (Google DeepMind)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Rapid retrieval-augmented generation and code reasoning Claude Haiku (Anthropic)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Lightweight LLM for fast inference and evaluation tasks GPT-5 Mini (OpenAI)     $\bullet$ Parameters: Not disclosed $|$ Architecture: Causal Decoder-Only Transformer     $\bullet$ Usage: Enhanced multi-turn reasoning and tool-augmented evaluation
TABLE VII: Evaluation of Qwen 2.5 7B, LLaMa 4 Maverick 17B, GPT-4o Mini, and Claude 3.5 Sonnet on PennyLane. Results for open-source models (Qwen and LLaMa) are highlighted in purple to distinguish their RAG-based performance gains from commercial models.
Metric Qwen 2.5 7B LLaMa 4 Maverick GPT-4o Mini Claude 3.5 Sonnet Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Full 75% 50% 0% Pass@1 33.0% 37.9% 33.3% 7.2% 67.8% 64.8% 67.8% 56.8% 54.5% 59.8% 56.4% 50.8% 70.8% 68.9% 64.4% 73.5% Pass@2 37.5% 41.3% 39.0% 8.0% 75.0% 76.5% 75.4% 67.4% 64.8% 71.6% 66.3% 66.3% 79.5% 82.6% 76.9% 87.1% Pass@3 39.4% 42.8% 41.3% 8.0% 80.3% 81.1% 79.2% 72.3% 72.0% 78.4% 73.9% 75.4% 84.5% 87.9% 81.1% 90.2% Pass@4 41.3% 44.3% 41.7% 8.3% 81.8% 82.2% 80.3% 76.5% 76.5% 82.9% 79.2% 81.4% 87.1% 91.7% 87.1% 93.2% Pass@5 41.7% 45.1% 43.2% 8.7% 84.8% 84.8% 82.2% 78.8% 80.3% 84.8% 81.8% 85.6% 89.0% 92.4% 87.5% 95.1%
Figure 6: Ablation Study: Vanilla RAG vs. Graph RAG evaluated on Gemini 2.5 Flash and Claude Haiku.
Analysis: Vanilla RAG vs. GraphRAG.
Figure 6 compares vanilla retrieval-augmented generation (RAG) against a GraphRAG variant on our PennyLane benchmark, using two representative commercial models (Gemini 2.5 Flash and Claude Haiku). Across all pass@k metrics, vanilla RAG consistently outperforms GraphRAG for both models, indicating that the added graph-structuring step does not translate into better executable-code success in this setting. Overall, these results suggest that, for our dataset and tasks, the primary gains come from retrieving a small set of directly relevant PennyLane exemplars rather than from explicitly modeling inter-document relations via a graph.
Since our benchmark emphasizes functional correctness of generated notebooks, where small syntactic or API-level deviations can cause execution failures, the precision of the retrieved snippets is often more important than broader relational coverage. Graph construction can also impose additional routing/selection constraints (e.g., traversing nodes that are only indirectly related), which may reduce the density of immediately usable code patterns in the final prompt compared to standard similarity-based retrieval. Given that vanilla RAG consistently outperforms GraphRAG in this ablation study, we adopt vanilla RAG as the default retrieval strategy for the remainder of the evaluations.
@@
Given that the dataset was primarily developed to support open-source development, our core evaluation focused on Qwen 7B and LLaMa 4. As shown in Table VII, these models showed clear improvements when used with our dataset. Notably, RAG with 75% context coverage often outperformed the full-context setting. This suggests that a moderate amount of relevant context may reduce noise and improve grounding, while excessive retrieval might introduce redundancy or irrelevant tokens.
@@
For comparison, we also tested GPT-4o Mini, Claude 3.5 Sonnet, Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini. As shown in Tables VII and VIII, these commercial models demonstrated high performance even without RAG. However, the added context from our dataset provided limited benefit. This may be attributed to their broad pretraining, which likely includes similar material (e.g., PennyLane documentation or equivalent quantum code examples), reducing the marginal gain from external retrieval.
TABLE VIII: Evaluation of Gemini 2.5 Flash, Claude Haiku, and GPT-5 Mini on PennyLane under two RAG coverage settings (75% and 0%).
Metric Gemini 2.5 Flash Claude Haiku GPT-5 Mini 75% 0% 75% 0% 75% 0% Pass@1 53.0% 55.3% 72.0% 67.4% 73.9% 74.6% Pass@2 69.3% 64.8% 81.8% 78.8% 85.6% 85.2% Pass@3 73.5% 68.9% 85.6% 84.1% 89.8% 88.3% Pass@4 79.2% 72.0% 88.3% 87.9% 90.5% 89.8% Pass@5 81.4% 75.0% 89.0% 89.8% 91.3% 92.0%